OpenStudy (anonymous):

Please please someone help!! :( A rock is launched from a cannon. It's height, h(x), can be represented by a quadratic function in terms of time, x, in seconds. After one second, the rock is 129 feet in the air; after 2 seconds, it is 236 feet in the air. Find the height, in feet, of the rock after 5 seconds in the air.

OpenStudy (matlee):

Do you know the formulas

OpenStudy (anonymous):

No I don't

OpenStudy (matlee):

Ok

OpenStudy (matlee):

I have a similar problem to this so we can just change the numbers

OpenStudy (anonymous):

Okay

OpenStudy (matlee):

A rock is launched from a cannon its height, h(x), can be represented by a quadratic function in terms of time, x, in seconds. After 1 second, the rock is 186 feet in the air; after 2 seconds, it is 344 in the air. Find the height, in feet,of the rock after 8 seconds in the air.; : Using the form: ax^2 + bx + c = y y = h(x), c = 0 : x=1; y=186 a(1^2) + 1b = 186 a + b = 186 : x=2; y=344 a(2^2) + 2b = 344 4a + 2b = 344 : Multiply the 1st equation by 2, subtract from the 2nd equation 4a + 2b = 344 2a + 2b = 372 --------------Subtraction eliminates b, find a 2a = -28 a = -14 : Find b using: a + b = 186 -14 + b = 186 b = 186 + 14 b = 200 : The equation: h(x) = -14x^2 + 200x : "Find the height, in feet,of the rock after 8 seconds in the air.; x=8 h(8) = -14(8^2) + 200(8) h(8) = -14(64) + 1600 h(8) = -896 + 1600 h(8) = 704 ft after 8 sec

OpenStudy (matlee):

Do you need help or just answer?

OpenStudy (anonymous):

I would prefer the answer, but help is ok. I tried to plug in all the numbers with what you gave me... but I just confuse myself. I really need to get this one right

OpenStudy (anonymous):

So if you could give me the answer, I'd be very thankful... and it'd be much appreciated!

OpenStudy (matlee):

Ok sure let me figure it out do you have time?

OpenStudy (anonymous):

Yes, take your time. Thank you so much!!

OpenStudy (matlee):

Your welcome!

OpenStudy (matlee):

315ft at 5 seconds

OpenStudy (matlee):

is the answer

OpenStudy (matlee):

Was it right?

OpenStudy (anonymous):

No I got it wrong, I just now have to get the last one right. It's no problem. Do you have time to maybe help me with one last one?

OpenStudy (matlee):

Sure

OpenStudy (anonymous):

Quick Computing Company produces calculators. They have found that the cost, c(x), of making x calculators is a quadratic function in terms of x. The company also discovered that it costs $45 to produce 2 calculators, $143 to produce 4 calculators, and $869 to produce 10 calculators. Find the total cost of producing 1 calculator.

OpenStudy (matlee):

Is there any options?

OpenStudy (matlee):

or are you typing?

OpenStudy (anonymous):

No options